# End-to-end transfer learning for speaker-independent cross-language and cross-corpus speech emotion recognition

### v2

@@ 4.1.1 Berlin Database of Emotional Speech (Emo-DB)
The Emo-DB [38] dataset is a German emotive speech dataset that has been widely used in SER research. This dataset consists of 535 utterances including 7 basic emotion catalogues (anger, boredom, disgust, anxiety/fear, happiness, sadness, and neutral). The utterances are performed by 10 professional actors, with 10 pre-defined sentences. Each sentence is performed with all different emotions, and the sentence content should not deliver sentimental information.
@@ 4.1.2 Ryerson Audio-Visual Database of Emotional Speech and Song (RAVDESS)
The RAVDESS [39] dataset contains recordings with 24 professional actors (12 female, 12 male), vocalising two lexically-matched statements (i.e. “Kids are talking by the door.”, “Dogs are sitting by the door.”). The speech recordings include calm, happy, sad, angry, fearful, surprise, and disgust expressions. Each expression is produced at two levels of emotional intensity (normal, strong), with an additional neutral expression, resulting in a total of 8 emotion catalogues.
@@ 4.1.3 Emotional Speech Dataset (ESD)
The ESD [40] dataset is a recent multilingual and multi-speaker emotional speech dataset designed for various speech synthesis and voice conversion tasks. The dataset consists of 350 parallel utterances spoken by 10 native English and 10 native Mandarin speakers. In this work, we only use the Mandarin utterances, which involve 5 male speakers and 5 female speakers, and are performed in 5 emotion catalogues (happy, sad, neutral, angry, and surprise).
@@
Reference [38] [39] [40] [40] [40]
@@ 4.2 Baseline features
For comparison with the proposed model, we choose the baseline systems using the most widely used feature sets designed for SER. The Extended Geneva Minimalistic Acoustic Parameter Set (eGeMAPS) [41] and the Emobase [42] feature sets produce features per utterance, computed from a set of low-level-descriptors to which various statistical functions are applied. Both the eGeMAPS and the Emobase feature sets contain spectral features (e.g., pitch-related, formant-related features), energy/amplitude features (e.g., loudness), and filter-bank features (e.g., MFCC). In Emobase, statistical functions including min/max, arithmetic mean, standard deviation, skewness, and kurtosis, are applied to all the low-level descriptors and their delta coefficients. In contrast, in eGeMAPS, only selected statistical functions are applied to some of the low-level descriptors, resulting a feature set containing a few extra temporal features such as the rate of loudness peaks features, and the mean length of voiced regions and unvoiced regions. The low-level descriptor selection criteria for eGeMAPS are based on the theoretical significance of the feature, the feature usage frequency in literature, and the potential of an acoustic parameter to indicate physiological changes in voice production during affective processes [41]. In term of the feature vector dimensionality, the Emobase feature vector has length 988, and the eGeMAPS feature vector has length 88.
@@ 4.3 Training on multiple languages
Similarly to [18, 19], we consider all combinations of merging two out of three language datasets for training and validation in order to meet the data requirements of the data-driven model and to allow the model to capture the variations across languages and corpora thus not to over-fit to one dataset.
@@ 4.5 Experimental settings
The Adagrad optimiser [43] is used to train the models, with a fixed learning rate of $3\times 10^{-4}$ and weight decay. The batch size is 14 due to memory constraints. The weighting parameter $\alpha$ in (3) is equal to 0.1 as in [26]. The weight decay parameter, the dropout rate, and the hyper-parameters $\beta$ and $\gamma$ in (7) and (9) are optimised using Hyperopt [44]. The optimal results are shown in Table 3 and correspond to the values that are used in the experiments.
@@ 5.4 Influence of training set size of target language
Rehman et. al. [45] Hybrid Neural Network (NN) EN+DE RAVDESS 4 - 56.2%
@@
Latif et. al. [22] Domain adaptation DE Emo-DB 5 - 81.3%
Parry et. al. [46] Datasets aggregation EN+IT+DE Emo-DB 3 - 69.72%
@@
Desplanques and Demuynck [16] Emotion i-vector DE Emo-DB 4 90.5% -
@@
Rehman et. al. [45] Hybrid NN EN+EN Emo-DB 4 - 40.2%
Zhang et. al. [21] Subspace learning EN Emo-DB 5 47.35% -
@@
Latif et. al. [22] Domain adaptation IT + UR + EN Emo-DB 5 - 68%
Parry et. al. [46] Datasets aggregation EN Emo-DB 3 - 41.99%
Desplanques and Demuynck [16] Emotion i-vector EN Emo-DB 4 81.4% -
@@
Table 4: Comparison to recent work in cross-language/cross-corpus SER. EN, DE, IT, CH, TU, and UR indicate English, German, Italian, Chinese, Turkish, and Urdu language, respectively.
@@ 5.5 Comparison to existing works
We also compare our method with a few recent methods proposed for cross-language/cross-corpus SER. The comparison is in Table 4 where we only present the results reported on the same testing datasets (i.e. on the Emo-DB and RAVDESS) as our work. Since there is hardly any research using ESD for cross-language SER, we do not include the results for this dataset.
@@ 6 Conclusions and future work
To alleviate the performance degradation in cross-language/cross-corpus SER compared to within-language/within-corpus SER, we proposed a transfer learning method that firstly uses the wav2vec 2.0 pre-trained model to transfer a time-domain audio waveform into a contextual embedding space that is shared across different languages, thereby reducing the language variabilities in the speech features. Then, by applying a Deep-WCCN layer, which is adapted to cope with DNN training, this Deep-WCCN layer can further reduce within-class variance caused by other factors (e.g. speaker identity, channel variability). Experimental results first show that the proposed method largely increases both within-language and cross-language SER performance compared to the eGeMAPS and Emobase feature sets that have been designed for and widely used in SER. Furthermore, an ablation study shows that Deep-WCCN can reduce the within-class variances which further improves the performance for the proposed DNN model. In contrast, the conventional WCCN does not show improvements on the eGeMAPS and Emobase feature sets with a RF classifier. Next, we show that the proposed transfer learning method exhibits good data efficiency in merging target language data in the fine-tuning process. The model speaker-independent SER performance increases for all testing target languages, and a performance gain in WA up to 15.6% and in UA up to 13.8% is achieved when only 160 s of target language data is used in the training set for fine-tuning. Finally, a comparison with recent work in cross-language/cross-corpus SER demonstrates that the proposed method can significantly improve within-language and cross-language SER performance among multiple datasets.
Future work firstly include the evaluation of the wav2vec 2.0 model pre-trained on an even larger speech dataset and the assessment of its benefit to cross-language SER performance. A good candidate dataset is proposed in [47] and consist of half a million hours of publicly available speech audio in 128 languages. Secondly, we plan to evaluate the proposed method on more emotive speech datasets, and to fine-tune the model on emotive speech datasets from more languages to create a general data-driven model for SER. Thirdly, a fine-tuning incorporating rank decomposition matrices as proposed in [48] could be tested under the scope of this paper. The fine-tuning scheme in [48] is expecting to retain the knowledge captured by the trained models, therefore improving the efficiency when adapting the model to more emotive speech datasets. Lastly, it may be relevant to reduce the model size through model distillation [49] to make it feasible for mobile and low-power devices.

### v3

@@ 4.1.1 Berlin Database of Emotional Speech (Emo-DB)
The Emo-DB [Burkhardt2005] dataset is a German emotive speech dataset that has been widely used in SER research. This dataset consists of 535 utterances including 7 basic emotion catalogues (anger, boredom, disgust, anxiety/fear, happiness, sadness, and neutral). The utterances are performed by 10 professional actors, with 10 pre-defined sentences. Each sentence is performed with all different emotions, and the sentence content should not deliver sentimental information.
@@ 4.1.2 Ryerson Audio-Visual Database of Emotional Speech and Song (RAVDESS)
The RAVDESS [Livingstone2018] dataset contains recordings with 24 professional actors (12 female, 12 male), vocalising two lexically-matched statements (i.e. “Kids are talking by the door.”, “Dogs are sitting by the door.”). The speech recordings include calm, happy, sad, angry, fearful, surprise, and disgust expressions. Each expression is produced at two levels of emotional intensity (normal, strong), with an additional neutral expression, resulting in a total of 8 emotion catalogues.
@@ 4.1.3 Emotional Speech Dataset (ESD)
The ESD [Zhou2021] dataset is a recent multilingual and multi-speaker emotional speech dataset designed for various speech synthesis and voice conversion tasks. The dataset consists of 350 parallel utterances spoken by 10 native English and 10 native Mandarin speakers. In this work, we only use the Mandarin utterances, which involve 5 male speakers and 5 female speakers, and are performed in 5 emotion catalogues (happy, sad, neutral, angry, and surprise).
@@
Reference [Burkhardt2005] [Livingstone2018] [Zhou2021] [Zhou2021] [Zhou2021]
@@ 4.2 Baseline features
For comparison with the proposed model, we choose the baseline systems using the most widely used feature sets designed for SER. The Extended Geneva Minimalistic Acoustic Parameter Set (eGeMAPS) [Eyben] and the Emobase [Eyben2009] feature sets produce features per utterance, computed from a set of low-level-descriptors to which various statistical functions are applied. Both the eGeMAPS and the Emobase feature sets contain spectral features (e.g., pitch-related, formant-related features), energy/amplitude features (e.g., loudness), and filter-bank features (e.g., MFCC). In Emobase, statistical functions including min/max, arithmetic mean, standard deviation, skewness, and kurtosis, are applied to all the low-level descriptors and their delta coefficients. In contrast, in eGeMAPS, only selected statistical functions are applied to some of the low-level descriptors, resulting a feature set containing a few extra temporal features such as the rate of loudness peaks features, and the mean length of voiced regions and unvoiced regions. The low-level descriptor selection criteria for eGeMAPS are based on the theoretical significance of the feature, the feature usage frequency in literature, and the potential of an acoustic parameter to indicate physiological changes in voice production during affective processes [Eyben]. In term of the feature vector dimensionality, the Emobase feature vector has length 988, and the eGeMAPS feature vector has length 88.
@@ 4.3 Training on multiple languages
Similarly to [Latif2018, Gerczuk2021], we consider all combinations of merging two out of three language datasets for training and validation in order to meet the data requirements of the data-driven model and to allow the model to capture the variations across languages and corpora thus not to over-fit to one dataset.
@@ 4.5 Experimental settings
The Adagrad optimiser [Duchi2010] is used to train the models, with a fixed learning rate of $3\times 10^{-4}$ and weight decay. The batch size is 14 due to memory constraints. The weighting parameter $\alpha$ in (3) is equal to 0.1 as in [Conneau2020]. The weight decay parameter, the dropout rate, and the hyper-parameters $\beta$ and $\gamma$ in (7) and (9) are optimised using Hyperopt [Bergstra2013]. The optimal results are shown in Table 3 and correspond to the values that are used in the experiments.
@@ 5.4 Influence of training set size of target language
Rehman et. al. [Rehman2020] Hybrid Neural Network (NN) EN+DE RAVDESS 4 - 56.2%
@@
Latif et. al. [Latif2019] Domain adaptation DE Emo-DB 5 - 81.3%
Parry et. al. [Parry2019] Datasets aggregation EN+IT+DE Emo-DB 3 - 69.72%
@@
Desplanques and Demuynck [Desplanques2018] Emotion i-vector DE Emo-DB 4 90.5% -
@@
Rehman et. al. [Rehman2020] Hybrid NN EN+EN Emo-DB 4 - 40.2%
Zhang et. al. [Zhang2020] Subspace learning EN Emo-DB 5 47.35% -
@@
Latif et. al. [Latif2019] Domain adaptation IT+UR+EN Emo-DB 5 - 68%
Parry et. al. [Parry2019] Datasets aggregation EN Emo-DB 3 - 41.99%
Desplanques and Demuynck [Desplanques2018] Emotion i-vector EN Emo-DB 4 81.4% -
Ma et. al. [ma2024emotion2vec] Self-supervised learning EN+CH+IT+ FR+BN+PS+ UR+GR+RU Emo-DB 4 84.34% 84.85%
@@
Table 4: Comparison to recent work in cross-language/cross-corpus SER. EN, DE, IT, CH, TU, FR, BN, PS, GR, RU and UR indicate English, German, Italian, Chinese, Turkish, French, Bengali, Persian, Greek, Russian and Urdu language, respectively.
@@ 5.5 Comparison to existing works
We also compare our method with a few recent methods proposed for speaker-independent cross-language/cross-corpus SER. The comparison is in Table 4 where we only present the results reported on the same testing datasets (i.e. on the Emo-DB and RAVDESS) as our work. Since there is hardly any research using ESD for cross-language SER, we do not include the results for this dataset.
@@ 6 Conclusions and future work
To alleviate the performance degradation in cross-language/cross-corpus SER compared to within-language/within-corpus SER, we proposed a transfer learning method that firstly uses the wav2vec 2.0 pre-trained model to transfer a time-domain audio waveform into a contextual embedding space that is shared across different languages, thereby reducing the language variabilities in the speech features. Then, by applying a Deep-WCCN layer, which is adapted to cope with DNN training, this Deep-WCCN layer can further reduce within-class variance caused by other factors (e.g. speaker identity, channel variability). Experimental results first show that the proposed method largely increases both within-language and cross-language SER performance compared to the eGeMAPS and Emobase feature sets that have been designed for and widely used in SER. Furthermore, an ablation study shows that Deep-WCCN can reduce the within-class variances which further improves the performance for the proposed DNN model. In contrast, the conventional WCCN does not show improvements on the eGeMAPS and Emobase feature sets with a RF classifier. Next, we show that the proposed transfer learning method exhibits good data efficiency in merging target language data in the fine-tuning process. The model speaker-independent SER performance increases for all testing target languages, and a performance gain in WA up to 15.6% and in UA up to 13.8% is achieved when only 160 s of target language data is used in the training set for fine-tuning. Finally, a comparison with recent work in cross-language/cross-corpus SER demonstrates that the proposed method can significantly improve speaker-independent within-language and cross-language SER performance among multiple datasets.
Future work firstly include the evaluation of the wav2vec 2.0 model pre-trained on an even larger speech dataset and the assessment of its benefit to cross-language SER performance. A good candidate dataset is proposed in [Babu2021] and consist of half a million hours of publicly available speech audio in 128 languages. Secondly, we plan to evaluate the proposed method on more emotive speech datasets, and to fine-tune the model on emotive speech datasets from more languages to create a general data-driven model for SER. Thirdly, a fine-tuning incorporating rank decomposition matrices as proposed in [hu2022lora] could be tested under the scope of this paper. The fine-tuning scheme in [hu2022lora] is expecting to retain the knowledge captured by the trained models, therefore improving the efficiency when adapting the model to more emotive speech datasets. Lastly, it may be relevant to reduce the model size through model distillation [Hinton2015] to make it feasible for mobile and low-power devices.